### v3

@@ Abstract
Context. All numerical solutions of the pulsar magnetosphere over the past 25 years show closed-line regions that end a significant distance inside the light cylinder, and manifest thick strongly dissipative separatrix surfaces instead of thin current sheets, with a tip that has a distinct pointed Y shape instead of a T shape. We need to understand the origin of these results which were not predicted by our early theories of the pulsar magnetosphere.
@@ 1 Gaps in our current understanding
Right after the discovery of pulsars (Hewish, 1968), Goldreich and Julian formulated the mathematical description of their plasma-filled magnetospheres in the ideal force-free limit in steady-state and axisymmetry (Goldreich & Julian, 1969). It took us three decades to produce the first solution of this idealized problem, mainly because of the mathematical singularity along the light cylinder (Contopoulos et al., 1999). Since then, this solution became a reference solution against which the results of numerical simulations are evaluated. As we will see below, the ideal force-free steady-state 3D problem is also well formulated mathematically, but still lacks a reference solution. This is mainly due to the presence of electric current sheets within the magnetosphere which strongly complicate all our standard numerical approaches.
@@
Recent advances in global particle-in-cell (PIC) simulations (e.g. Philippov et al., 2015; Hu & Beloborodov, 2022; Hakobyan et al., 2023; Bransgrove et al., 2023; Soudais et al., 2024) yield separatrix regions between open and closed field lines that exhibit a significant thickness beyond the simulation resolution (see Appendix), and a multi-layered internal structure. This is intriguing because, the same simulations are able to generate Harris-type equatorial current sheets with thickness at the limit of their numerical resolution (the skin-depth). Moreover, their closed-line regions terminate well inside the light cylinder, with regions of strong electromagnetic dissipation beyond their tips, something that was not expected by Goldreich and Julian. Interestingly, the shape of the tips resembles a pointed Y, thus deviating from the theoretically predicted T shape (Uzdensky, 2003). Furthermore, the extent of the closed-line region and the sharpness of the Y appears to be linked to the efficiency of pair production in the simulation (see comment in Kalapotharakos et al., 2018). Finally, our recent work in Contopoulos et al. (2024) indicates that as the closed-line region approaches the light cylinder, the electric current in the separatrix decreases, and in the limit where it touches the light cylinder, the equatorial current sheet vanishes entirely.
@@
The origin of these results remains unclear. Some researchers suspect that they arise either due to insufficient resolution in current PIC simulations, or to inadequate running times to reach a true steady-state. On the other hand, the resolution of current PIC simulations falls significantly short in modeling the microphysics of the equatorial current sheet. To generate pulsar light curves and spectra for comparison with observations, simulation results are extrapolated by several orders of magnitude (particle Lorentz factors, magnetic field values, electromagnetic spectra, etc.). Unfortunately, there is no consensus among research groups regarding the specific extrapolation method. Consequently, a universally accepted understanding of the physical origin of high-energy radiation from pulsars remains elusive.
@@ 2 Domain decomposition
×$\nabla\times\left\{\left(1-\left[\frac{r\sin\theta}{R_{\rm lc}}\right]^{2}\right){\bf B}_{\rm p}+{\bf B}_{\phi}\right\}-\alpha{\bf B}=0\ ,$ (3)
@@
The new methodology, first proposed in Paper I to avoid the mathematical problems associated with the magnetospheric current sheet, is the domain decomposition into open and closed field lines (see schematic in figure 1 of that paper). In particular, we choose from the very beginning of our simulation which field lines will be open and which ones will be closed. This is equivalent to an ad hoc determination from the very beginning of the extent of the polar cap (see below). In general, the solution that we will obtain will contain a closed-line region that does not extend all the way to the light cylinder. Obviously, the choice of the extent and shape of the polar cap is arbitrary, thus the solution is degenerate, as in the axisymmetric analyses of Contopoulos (2005) and Timokhin (2006). [^1] In the early literature (i.e. before the era of ab-initio global PIC simulations; Philippov et al., 2015; Kalapotharakos et al., 2018) it was expected that the single solution that nature will choose is the one in which the closed-line region has its maximum extent and everywhere touches the light cylinder. We have seen, however, in the PIC literature of the past 10 years that this is not the case and that the closed-line region ends some significant distance inside the light cylinder. Contopoulos et al. (2024) offered a tentative explanation for this effect. The issue of the extent of the closed-line region is further complicated after Paper II who concluded that, when the closed-line region reaches the light cylinder, the whole magnetosphere is contained inside the light cylinder, the open-line region disappears altogether, and the pulsar dies out completely. Several numerical simulations exist in the literature in which the closed-line region touches the light cylinder (e.g. Contopoulos et al., 1999; Timokhin, 2006; Spitkovsky, 2006, etc.), thus we need to understand the origin of this discrepancy.
Once the size and shape of the polar cap is determined, the shape of the separatrix surface is unknown and must be self-consistently determined. In the present paper, we implement in 3D the procedure first proposed in Paper I. We begin by separating the regions of open and closed field lines with an ad hoc chosen surface, namely
@@
that originates at the edges of the polar caps around the magnetic poles of the central star. A natural (but certainly not unique) initial choice might be a dipolar surface that ends on a circular polar cap centered around the magnetic axis. The reason we decided to separate the two regions is that, due to the electric current sheet that flows along the separatrix between the two regions, a strong contact discontinuity (of practically infinitesimal thickness) in the magnetic and electric fields develops along the separatrix. Uzdensky (2003) integrated eq. (2) across current sheets and obtained that
@@
    Solve eqs. (3) and (4) independently in the two regions for an initial arbitrary choice of the separatrix surface between them.
@@
    Implement the condition that the separatrix surface lies along magnetic field lines, namely $\hat{\bf n}\cdot{\bf B}=0$ right above and below the separatrix surface, where $\hat{\bf n}$ is a vector perpendicular to the separatrix surface.
@@
    Adjust the shape of the separatrix so as to gradually remove the discontinuity of $B^{2}-E^{2}$ over its surface ($B^{2}-E^{2}$ will in general be discontinuous for an ad hoc choice of the separatrix; see Papers I & II for details).
@@
    Repeat all of the above steps after each re-adjustment of the separatrix till pressure balance is satisfied at all points and the separatrix does not need to re-adjust its shape anymore. The final steady-state solution will thus be obtained.
@@
Figure 1: Evolution of the total NN losses for the solution with $\lambda=20^{\circ}$ and $\theta_{\rm pc}=36^{\circ}$ in the closed-line (blue) and open-line (orange) regions with respect to the number of training epochs (the difference in the number of epochs has to do with the different number of steps the training algorithm takes during its second-order optimization). We started the training with a first order optimizer and switched to a second order one after $10^{4}$ training epochs. Beyond that, we adjusted the shape of the separatrix every $3\times 10^{4}$ epochs. The achieved total training loss between $10^{-5}$ and $10^{-6}$ is deemed satisfactory.
Figure 2: Cross section of steady-state solution representing an inclined rotator with $\lambda=20^{\circ}$ and $\theta_{\rm pc}=36^{\circ}$. In this simulation $r_{*}=0.25R_{\rm lc}$. Rotation axis along $z$. The inclined magnetic axis lies along the corotating $xz$ plane shown. This particular cross section represents phase 00 (or $\pi$) of the corresponding rotating time-dependent solution. Closed thick black lines: separatrix between open and closed field lines. Open thick black lines: separatrix between open field lines that originate from the north and south magnetic poles. This is where the equatorial current sheet lies. Red lines: initial dipolar shapes of the separatrices before readjustment. For this particular choice of the polar cap, the dipole is significantly stretched outwards closer to the light cylinder (represented by the two dashed lines at $x/R_{\rm lc}=\pm 1$). The equatorial current sheet lies where lines from the north and south polar cap rim meet. Color scale: ratio $B_{p}/B$. This represents the development of the azimuthal magnetic field $B_{\phi}$ accross the magnetosphere. Notice that at the magnetospheric Y-point where the equatorial current sheet connects to the separatrix current sheet, $B_{p}=0$ and $B_{\phi}\neq 0$ as expected (Uzdensky, 2003).
Figure 3: Same as Fig. 2 but along the corotating $yz$ plane. This particular cross section represents phase $\pi/2$ (or $23\pi/$ 2) of the corresponding rotating time-dependent solution.
Figure 4: The steady-state solution of Fig. 2 in 3D as seen from above. Yellow cylinder: light cylinder. Pale undulating surface: equatorial current sheet. We plot only open magnetic field lines just inside the rim of the northern polar cap. We see the clear azimuthal break of open field lines expected very close to the Y-point where $B_{p}\rightarrow 0$ and $B_{\phi}\neq 0$.
@@ 3 Results
We experimented with various NN configurations, number layers and nodes per layer, various optimization procedures, and various activation functions. This tuning of hyperparameters is a tedious manual process that requires significant effort before the first meaningful results are obtained. In the solution presented below we implemented two Fully Connected NNs with 3 entires ($x,y,z$), 5 internal layers consisting of 128 nodes each, and 4 exits ($B_{x},B_{y},B_{z},\alpha$), first order Adam optimizers initially and second order Broyden-Fletcher-Goldfarb-Shanno (BFGS) optimizers later on, tanh activation functions, and roughly 270,000 training points distributed randomly over the various magnetospheric computational domains and their boundaries. Notice that, in order to avoid overfitting, the training points are periodically updated. The reason we implemented a cartesian system of coordinates was to avoid the singularities of the spherical system of coordinates around the axis $\theta=0$ and the requirement for periodicity between $\phi=0$ and $\phi=2\pi$. We validated our procedure by training our NNs with the analytic vacuum dipole solution of ×$\nabla\times{\bf B}=0$ (no rotation, no light cylinder, no electric currents). The comparison with the analytic dipolar solution was fair (see also Paper I). This gave us confidence to address the full 3D pulsar magnetosphere problem with PINNs. In Fig. 1 we show the evolution of the total training losses for our main NNs. The shape of the separatrix was re-adjusted every 30,000 training epochs. We performed 25 readjustments (high spikes in Fig. 1) in order to reach a solution with individual requirements satisfied to a precision of a few times $10^{-6}$ in both NNs. The solution relaxes to a particular separatrix shape that satisfies pressure balance at all points accross it. The training of our two NNs required about 24 hours of computational time on a standard off-the-shelf computer with a GPU. We thus obtained a dissipationless ideal steady-state solution for a star with radius $r_{*}=0.25R_{\rm lc}$, a uniform polar cap with constant angular opening $\theta_{\rm pc}=36^{\circ}(\pi/5)$ around the magnetic axis, and an inclination angle of $\lambda=20^{\circ}(\pi/9)$. As we will see next, our first results are promising and emphasize the potential of our method.
@@ 4 Summary and conclusions
We have developed a novel robust numerical method that allows us to calculate steady-state force-free ideal 3D pulsar magnetospheres for various pulsar inclinations and polar cap shapes. Our method does not involve a time-dependent simulation that eventually relaxes to a steady-state, as in all previous attempts to obtain the steady-state pulsar magnetosphere. Instead, we solve directly the generalization of the axisymmetric pulsar equation (Scharlemann & Wagoner, 1973) in 3D (Endean, 1974; Mestel, 1975). Notice that the method is not limited to centrally located dipole fields and can accomodate any stellar magnetic field configuration. For example, one may determine a particular magnetic field configuration with hotspots that reproduce the NICER X-ray observations, assume that these hotspots correspond to the origin of open magnetic field lines (i.e. assume that these are the polar caps), and calculate the 3D ideal magnetosphere that corresponds to these particular polar caps using our methodology.
@@
In this paper, we have trained a set of two NNs that instantaneously yields the three components $B_{x},B_{y},B_{z}$ of the magnetic field and their spatial derivatives at any given point $(x,y,z)$ of the magnetosphere for one particular pulsar inclination, and one particular polar cap size and shape. This is extremely helpful if one wants to draw magnetospheric features like magnetic and electric field lines, electric charge and electric current distributions, Poynting flux, current sheets, etc. We will present a detailed investigation of the pulsar magnetosphere for various pulsar inclinations and for various polar cap sizes and shapes in an extended forthcoming publication, and we will produce corresponding trained NNs (Paper IV). We look forward to learn whether magnetospheric solutions exist even for arbitrarily small polar caps with arbitrarily small luminosities as seen in 2D (Paper II). We would like to emphasize that our methodology allows us to obtain new solutions of the ideal 3D pulsar magnetosphere by determining a priori various shapes and sizes of the polar cap. Such solutions were never obtained before because all numerical solutions obtained in the literature always relax to one particular magnetospheric configuration, i.e. to one particular polar cap shape and size determined by the simulation, without any possibility for control by the programmer. This opens up a whole spectrum of solutions between which the real pulsar magnetosphere may choose to transition during its evolution. More work is needed along these lines.

### v4

@@ Abstract
Context. All PIC simulations of the pulsar magnetosphere over the past decade show closed-line regions that end a significant distance inside the light cylinder, and manifest thick strongly dissipative separatrix surfaces instead of thin current sheets, with a tip that has a distinct pointed Y shape instead of a T shape. We need to understand the origin of these results which were not predicted by our earlier numerical simulations of the pulsar magnetosphere.
@@ 1 Gaps in our current understanding
Right after the discovery of pulsars (Hewish, 1968), Goldreich and Julian formulated the mathematical description of their plasma-filled magnetospheres in the ideal force-free limit in steady-state and axisymmetry (Goldreich & Julian, 1969). It took us three decades to produce the first solution of this idealized problem, mainly because of the mathematical singularity along the light cylinder (Contopoulos et al., 1999). Since then, this solution became a reference solution against which the results of numerical simulations are evaluated. As we will see below, the ideal force-free steady-state 3D problem is also well formulated mathematically, but still lacks a reference solution. This is mainly due to the presence of electric current sheets within the magnetosphere which strongly complicate all our standard numerical approaches [^1] In axisymmetry, the current sheet may be fixed to lie along the boundary of the computational domain in the equator as in Contopoulos et al. (1999), while in 3D the position of the current sheet must be determined by the simulation. If one decides to simulate the full axisymmetric magnetosphere as in Cerutti et al. (2015), the same complication arises since the current sheet is free to oscillate around the equator..
@@
Recent advances in global particle-in-cell (PIC) simulations (e.g. Philippov et al., 2015; Hu & Beloborodov, 2022; Hakobyan et al., 2023; Bransgrove et al., 2023; Soudais et al., 2024) yield separatrix regions between open and closed field lines that exhibit a significant thickness beyond the simulation resolution (see Appendix), and a multi-layered internal structure. This is intriguing because, the same simulations are able to generate Harris-type equatorial current sheets with thickness at the limit of their numerical resolution (the skin-depth). Moreover, their closed-line regions terminate well inside the light cylinder, with regions of strong electromagnetic dissipation beyond their tips, something that was not expected by Goldreich and Julian. Interestingly, the shape of the tips resembles a pointed Y, thus deviating from the theoretically predicted T shape (Uzdensky, 2003). Furthermore, the extent of the closed-line region and the sharpness of the Y appears to be linked to the efficiency of pair production in the simulation (see comment in Kalapotharakos et al., 2018). Finally, our recent work in Contopoulos et al. (2024) indicates that as the closed-line region approaches the light cylinder, the electric current in the decreases, and in the limit where it touches the light cylinder, the equatorial current sheet vanishes entirely.
@@
The origin of these results remains unclear. Some researchers suspect that they arise either due to insufficient resolution in current PIC simulations, or to inadequate titletimes to reach a true steady-state. On the other hand, the resolution of current PIC simulations falls significantly short in modeling the microphysics of the equatorial current sheet. To generate pulsar light curves and spectra for comparison with observations, simulation results are extrapolated by several orders of magnitude (particle Lorentz factors, magnetic field values, electromagnetic spectra, etc.). Unfortunately, there is no consensus among research groups regarding the specific extrapolation method. Consequently, a universally accepted understanding of the physical origin of high-energy radiation from pulsars remains elusive.
@@ 2 Domain decomposition
$\nabla\times\left\{\left(1-\left[\frac{r\sin\theta}{R_{\rm lc}}\right]^{2}\right){\bf B}_{\rm p}+{\bf B}_{\phi}\right\}-\alpha{\bf B}=0\ ,$ (3)
@@
The new methodology, first proposed in Paper I to avoid the mathematical problems associated with the magnetospheric current sheet, is the domain decomposition into open and closed field lines (see schematic in figure 1 of that paper). In particular, we choose from the very beginning of our simulation which field lines will be open and which ones will be closed. This is equivalent to an ad hoc determination from the very beginning of the extent of the polar cap (see below). In general, the solution that we will obtain will contain a closed-line region that does not extend all the way to the light cylinder. Obviously, the choice of the extent and shape of the polar cap is arbitrary, thus the solution is degenerate, as in the axisymmetric analyses of Contopoulos (2005) and Timokhin (2006).
In the early literature (i.e. before the era of ab-initio global PIC simulations; Philippov et al., 2015; Kalapotharakos et al., 2018) it was expected that the single solution that nature will choose is the one in which the closed-line region has its maximum extent and everywhere touches the light cylinder. We have seen, however, in the PIC literature of the past 10 years that this is not the case and that the closed-line region ends some significant distance inside the light cylinder. Contopoulos et al. (2024) offered a tentative explanation for this effect. The issue of the extent of the closed-line region is further complicated after Paper II who concluded that, when the closed-line region reaches the light cylinder, the whole magnetosphere is contained inside the light cylinder, the open-line region disappears altogether, and the pulsar dies out completely. Several numerical simulations exist in the literature in which the closed-line region touches the light cylinder (e.g. Contopoulos et al., 1999; Timokhin, 2006; Spitkovsky, 2006, etc.), thus we need to understand the origin of this discrepancy.
Once the size and shape of the polar cap are determined, the separatrix surface is also determined. Initially, however, its shape is unknown to us. We then implement in 3D the procedure first proposed in Paper I. We begin by separating the regions of open and closed field lines with an ad hoc chosen surface, namely
@@
that originates at the edges of the polar caps around the magnetic poles of the central star. We will denote this surface the ‘interim separatrix’. As we will see below, we will iteratively adjust its shape above the polar cap till it reaches the final shape of the true separatrix that corresponds to the fixed size and shape of the polar cap that we have chosen. A natural (but certainly not unique) initial choice might be a dipolar surface that ends on a circular polar cap centered around the magnetic axis. The reason we decided to separate the two regions is that, due to the electric current sheet that flows along the separatrix between the two regions, a strong contact discontinuity (of practically infinitesimal thickness) in the magnetic and electric fields develops along the separatrix. Uzdensky (2003) integrated eq. (2) across current sheets and obtained that
@@
    Solve eqs. (3) and (4) independently in the two regions for an initial arbitrary choice of the interim separatrix between them.
@@
    Implement the condition that the interim separatrix lies along magnetic field lines, namely $\hat{\bf n}\cdot{\bf B}=0$ right above and below it, where $\hat{\bf n}$ is a vector perpendicular to the interim separatrix.
@@
    Adjust the shape of the interim separatrix so as to gradually remove the discontinuity of $B^{2}-E^{2}$ over its surface ($B^{2}-E^{2}$ will in general be discontinuous for an ad hoc choice of the interim separatrix; see Papers I & II for details).
@@
    Repeat all of the above steps after each re-adjustment of the interim separatrix till pressure balance is satisfied at all points and the interim separatrix does not need to re-adjust its shape anymore. The true separatrix and the final steady-state solution will thus be obtained.
@@
Figure 1: Evolution of the total NN losses for the solution with $\lambda=20^{\circ}$ and $\theta_{\rm pc}=36^{\circ}$ in the closed-line (blue) and open-line (orange) regions with respect to the number of training epochs (the difference in the number of epochs has to do with the different number of steps the training algorithm takes during its second-order optimization). We started the training with a first order optimizer and switched to a second order one after $10^{4}$ training epochs. Beyond that, we adjusted the shape of the interim separatrix every $3\times 10^{4}$ epochs (blue and orange high spikes). The achieved total training loss between $10^{-5}$ and $10^{-6}$ is deemed satisfactory.
Figure 2: Cross section of steady-state solution representing an inclined rotator with $\lambda=20^{\circ}$ and $\theta_{\rm pc}=36^{\circ}$. In this simulation $r_{*}=0.25R_{\rm lc}$. Rotation axis along $z$. The inclined magnetic axis lies along the corotating $xz$ plane shown. This particular cross section represents phase 0 (or $\pi$) of the corresponding rotating time-dependent solution. Closed thick black lines: separatrix between open and closed field lines. Open thick black lines: separatrix between open field lines that originate from the north and south magnetic poles. This is where the equatorial current sheet lies. Red lines: initial dipolar shapes of the interim separatrices before readjustment. For this particular choice of the polar cap, the dipole is significantly stretched outwards closer to the light cylinder (represented by the two dashed lines at $x/R_{\rm lc}=\pm 1$). The equatorial current sheet lies where lines from the north and south polar cap rim meet. Color scale: ratio $B_{p}/B$. This represents the development of the azimuthal magnetic field $B_{\phi}$ accross the magnetosphere. Notice that at the magnetospheric Y-point where the equatorial current sheet connects to the separatrix current sheet, $B_{p}=0$ and $B_{\phi}\neq 0$ as expected (Uzdensky, 2003). Notice also that in this solution the Y-point is still significantly inside the light cylinder.
Figure 3: Same as Fig. 2 but along the corotating yzyz plane. This particular cross section represents phase $\pi/2$ (or $3\pi/2$) of the corresponding rotating time-dependent solution.
Figure 4: The steady-state solution of Fig. 2 in 3D as seen from above. Yellow cylinder: light cylinder. Pale undulating surface: equatorial current sheet. We plot only open magnetic field lines just inside the rim of the northern polar cap. The Y-point in this solution is significantly inside the light cylinder. We see the clear azimuthal break of open field lines expected very close to the Y-point where $B_{p}\rightarrow 0$ and $B_{\phi}\neq 0$.
@@ 3 Results
We experimented with various NN configurations, number layers and nodes per layer, various optimization procedures, and various activation functions. This tuning of hyperparameters is a tedious manual process that requires significant effort before the first meaningful results are obtained. In the solution presented below we implemented two Fully Connected NNs with 3 entires ($x,y,z$), 5 internal layers consisting of 128 nodes each, and 4 exits ($B_{x},B_{y},B_{z},\alpha$), first order Adam optimizers initially and second order Broyden-Fletcher-Goldfarb-Shanno (BFGS) optimizers later on, tanh activation functions, and roughly 270,000 training points distributed randomly over the various magnetospheric computational domains and their boundaries. Notice that, in order to avoid overfitting, the training points are periodically updated. The reason we implemented a cartesian system of coordinates was to avoid the singularities of the spherical system of coordinates around the axis $\theta=0$ and the requirement for periodicity between $\phi=0$ and $\phi=2\pi$. We validated our procedure by training our NNs with the analytic vacuum dipole solution of $\nabla\times{\bf B}=0$ (no rotation, no light cylinder, no electric currents). The comparison with the analytic dipolar solution was fair (see also Paper I). This gave us confidence to address the full 3D pulsar magnetosphere problem with PINNs. In Fig. 1 we show the evolution of the total training losses for our main NNs. The shape of the interim separatrix was re-adjusted every 30,000 training epochs. We performed 25 readjustments (high spikes in Fig. 1) in order to reach a solution with individual requirements satisfied to a precision of a few times $10^{-6}$ in both NNs. The solution relaxes to a particular separatrix shape that satisfies pressure balance at all points accross it. The training of our two NNs required about 24 hours of computational time on a standard off-the-shelf computer with a GPU. We thus obtained a dissipationless ideal steady-state solution for a star with radius $r_{*}=0.25R_{\rm lc}$, a uniform polar cap with constant angular opening $\theta_{\rm pc}=36^{\circ}(\pi/5)$ around the magnetic axis, and an inclination angle of $\lambda=20^{\circ}(\pi/9)$. As we will see next, our first results are promising and emphasize the potential of our method.
@@ 4 Summary and conclusions
We have developed a novel robust numerical method that allows us to calculate steady-state force-free ideal 3D pulsar magnetospheres for various pulsar inclinations and polar cap shapes. Our method does not involve a time-dependent simulation that eventually relaxes to a steady-state, as in all previous attempts to obtain the steady-state pulsar magnetosphere. Instead, we solve directly the generalization of the axisymmetric pulsar equation (Scharlemann & Wagoner, 1973) in 3D (Endean, 1974; Mestel, 1975). Notice that the method is not limited to centrally located dipole fields and can accomodate more complex stellar magnetic field configurations.
For example, one may determine a particular magnetic field configuration with hotspots that reproduce the NICER X-ray observations, assume that these hotspots correspond to the origin of open magnetic field lines (i.e. assume that these are the polar caps), and calculate the 3D ideal magnetosphere that corresponds to these particular polar caps using our methodology. The initial field configuration that would reproduce these polar cap shapes will require inclusion of higher multipoles. Very small polar caps can be produced by an aligned and displaced dipole and quadrupole (e.g. Gralla et al., 2017), and NICER hot spots require more complicated field structures with off-set dipole and quadrupole configurations (Chen et al., 2020; Kalapotharakos et al., 2021). This will complicate our choice of the initial ad hoc shape of the interim separatrix, but as long as the polar caps remain fixed, the iteratively adjustable shape of the interim separatrix will remain topologically the same, thus we believe that our procedure is directly applicable and will indeed eventually relax to the true shape of the separatrix.
@@
We would like to emphasize that the general steady-state ideal force-free solution is degenerate in the same sense that the axisymmetric solutions of Contopoulos (2005) and Timokhin (2006) are degenerate. In the axisymmetric case, the polar cap is circular, but its radial extent is arbitrary and depends on how far the closed-line region extends from the light cylinder (the larger the polar cap, the shorter the closed-line region). Obviously, the solution where the closed-line region touches upon the light cylinder is unique, and consequently, the corresponding extent of the polar cap is also uniquely defined. In our present 3D case, not only is the size of the polar cap arbitrary, but also its shape and position. The solution that we obtained in this work corresponds to a circular polar cap centered around the magnetic axis. This choice causes the closed field line boundary to lie at different distances with respect to the light cylinder as a function of azimuth and therefore does not capture correctly the field line sweep back at the polar cap. With our chosen inclination of $20^{\circ}$, this does not make much difference in the solution, but for larger inclination angles, it will make a big difference. We are certainly not claiming that this is the only ideal steady-state force-free solution of the pulsar magnetosphere. The general solution is clearly degenerate, and what we are presenting is just one valid solution. If we insist that the closed-line region everywhere touches upon the light cylinder, that solution will be unique and it will correspond to one unique polar cap shape that will not be circular nor centered around the magnetic axis. In fact, it has been shown in multiple studies that the polar caps on to which previous time-dependent numerical simulations converge are very different from those of a static (or even a retarded) vacuum dipole and depend on the magnetic inclination angle (e.g. Dyks & Harding, 2004; Timokhin & Arons, 2013).
In this paper, we have trained a set of two NNs that instantaneously yields the three components $B_{x},B_{y},B_{z}$ of the magnetic field and their spatial derivatives at any given point $(x,y,z)$ of the magnetosphere for one particular pulsar inclination, and one particular polar cap size and shape. This is extremely helpful if one wants to draw magnetospheric features like magnetic and electric field lines, electric charge and electric current distributions, Poynting flux, current sheets, etc. We will present a detailed investigation of the pulsar magnetosphere for various pulsar inclinations and for various polar cap sizes and shapes in an extended forthcoming publication, and we will produce corresponding trained NNs (Paper IV). We also plan to investigate the unique solution where the closed-line region everywhere touches upon the light cylinder. Determining this unique solution will require adjusting both the polar caps as well as the separatrix at every iteration step. We look forward to learn whether magnetospheric solutions exist even for arbitrarily small polar caps with arbitrarily small luminosities as seen in 2D (Paper II). We would like to emphasize that our methodology allows us to obtain new solutions of the ideal 3D pulsar magnetosphere by determining a priori various shapes and sizes of the polar cap. Such solutions were never obtained before because all numerical solutions obtained in the literature always relax to one particular magnetospheric configuration, i.e. to one particular polar cap shape and size determined by the simulation, without any possibility for control by the programmer. This opens up a whole spectrum of solutions between which the real pulsar magnetosphere may choose to transition during its evolution. More work is needed along these lines.